OpenStudy (anonymous):

One robot assembles all the needed parts in 8 hours, and another assembles them in 11 hours. How long will it take with both robots working together? Round to the nearest tenth.(Points : 1) 4.2 hours 2.2 hours 4.6 hours 3.2 hours

OpenStudy (anonymous):

help help help help? :DDD <3

OpenStudy (anonymous):

thnx jazy.:D!

OpenStudy (anonymous):

Your welcome! I took it off cuz the 19/88 thing was messed up, but either way the result is still 4.6:)

OpenStudy (anonymous):

Robot 1 does 1/8 of the job per hour Robot 2 does 1/11of the job per hour t = time working to finish job so 1/8t + 1/11t = 1 Solve for t. 1/88 mutliplied by 88: so 11t + 8t = 88 solve for t: t = 4.6 So it takes the 4.6hrs